# Protecting Federated Learning from Extreme Model Poisoning Attacks via Multidimensional Time Series Anomaly Detection

### v2

@@ 7 CONCLUSION
In the future, we will address the primary limitations of this work, as discussed in Appendix K.
@@
The (anonymous) GitHub repository with the code and the data to replicate the results discussed in this work is accessible at the following link: https://anonymous.4open.science/r/flanders_exp-7EEB/
Appendix A A Note on the Terminology Used
@@
Appendix B The Impact of Malicious Clients at each FL Round
@@
Appendix C The Intuition behind FLANDERS: Predictability of Legitimate vs. Malicious Local Models
In this section, we elaborate further on our findings discussed in Section 4.2.
We begin by considering again two clients $i,j\in\mathcal{C}$, where client $i$ is assumed to be a legitimate participant, while client $j$ acts maliciously. We assume that both clients are selected by the server over a sequence of consecutive $T$ FL rounds to train a global model on the MNIST dataset. Therefore, we examine the local models sent to the server at each round $t\in\{1,2,\ldots,T\}$ by clients $i$ and $j$. Specifically, these are $d$-dimensional vectors of real-valued parameters $\bm{\theta}_{i}^{(t)},\bm{\theta}_{j}^{(t)}\in\mathbb{R}^{d}$, such that $\bm{\theta}_{i}^{(t)}=(\theta_{i,1}^{(t)},\ldots,\theta_{i,d}^{(t)})$ and $\bm{\theta}_{j}^{(t)}=(\theta_{j,1}^{(t)},\ldots,\theta_{j,d}^{(t)})$, respectively.
To mimic the behavior of a hypothetical optimal FLANDERS filter, and therefore avoid the propagation of poisoned local models sent by client $j$ across the $T$ rounds, we assume that, at each round $t$, the global model from which both clients start their local training process is polished from any malicious updates received at the previous round $t-1$. Next, we calculate the average time-delayed mutual information (TDMI) for each pair of observed local models $(\bm{\theta}_{i}^{(t)},\bm{\theta}_{i}^{(t^{\prime})})$ and $(\bm{\theta}_{j}^{(t)},\bm{\theta}_{j}^{(t^{\prime})})$ across $T=50$ rounds, where $t^{\prime}>t$, for both client $i$ and $j$.
Firstly, we consider the special case where $t^{\prime}=t+1$ and compute the average TDMI between each pair of consecutive local models sent by the legitimate client and the malicious client, when this runs one of the four attacks considered in this work, namely GAUSS, LIE, OPT, and AGR-MM. In Figure 1, we plot the empirical distributions of the observed average TDMI for the legitimate and malicious clients.
[…]
Table 4: One-tailed $t$-test against the null hypothesis $H_{0}:\bar{\bm{\theta}_{i}}=\bar{\bm{\theta}_{j}}^{atk}$. Each cell contains the $p$-value for the statistical test corresponding to a specific attack. In all four cases, there is enough evidence to reject the null hypothesis at a confidence level $\alpha=0.01$ ($p$-value $\ll 0.01$).
$atk$
GAUSS LIE OPT AGR-MM
$H_{0}:\bar{\bm{\theta}_{i}}=\bar{\bm{\theta}_{j}}^{atk}$ $5.67*10^{-38}$ $8.33*10^{-5}$ $6.70*10^{-18}$ $4.20*10^{-12}$
Appendix D Matrix Autoregressive Model (MAR)
@@
$\begin{split}\ell(\bm{\widetilde{\Omega}};{\bm{\bm{\Theta}}}_{t})=||{\bm{\bm{\Theta}}}_{t}-\widetilde{{\bm{\bm{\Theta}}}}_{t}||^{2}_{\text{F}}=||{\bm{\bm{\Theta}}}_{t}-f(\bm{\Theta}_{t-1};\bm{\widetilde{\Omega}})||^{2}_{\text{F}}=||{\bm{\bm{\Theta}}}_{t}-\widetilde{{\bm{A}}}{\bm{\bm{\Theta}}}_{t-1}\widetilde{{\bm{B}}}||^{2}_{\text{F}}=\ell(\widetilde{{\bm{A}}},\widetilde{{\bm{B}}};{\bm{\bm{\Theta}}}_{t}),\end{split}$ (6)
@@
[…]
@@
$\frac{\partial{\mathcal{L}({\bm{A}},{\bm{B}};{\bm{\bm{\Theta}}}_{t},l)}}{\partial{{\bm{A}}}}$ $=\frac{\partial{\sum_{j=0}^{l-1}||{\bm{\bm{\Theta}}}_{t-j}-{\bm{A}}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}||^{2}_{\text{F}}}}{\partial{{\bm{A}}}}=$
$=-2\sum_{j=0}^{l-1}({\bm{\bm{\Theta}}}_{t-j}-{\bm{A}}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}){\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-1}=$
$=-2\Bigg{[}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}-{\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\Bigg{]}.$ (10)
@@
$-2\Bigg{[}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}-{\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\Bigg{]}=0$
@@
$\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}-{\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}=0.$
@@
$\begin{split}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}-{\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}=0\end{split}$ (11)
$\begin{split}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}={\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\end{split}$ (12)
$\begin{split}{\bm{A}}=\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}^{-1}.\end{split}$ (13)
@@
$\begin{split}{\bm{B}}=\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}^{T}_{t-j}{\bm{A}}{\bm{\bm{\Theta}}}_{t-j-1}\Bigg{)}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}^{T}_{t-j-1}{\bm{A}}^{T}{\bm{A}}{\bm{\bm{\Theta}}}_{t-j-1}\Bigg{)}^{-1}.\end{split}$ (15)
@@
${\bm{A}}^{(i+1)}=\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}^{(i)}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}({\bm{B}}^{(i)})^{T}{\bm{B}}^{(i)}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}^{-1};$
${\bm{B}}^{(i+1)}=\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}^{T}_{t-j}{\bm{A}}^{(i+1)}{\bm{\bm{\Theta}}}_{t-j-1}\Bigg{)}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}^{T}_{t-j-1}({\bm{A}}^{(i+1)})^{T}{\bm{A}}^{(i+1)}{\bm{\bm{\Theta}}}_{t-j-1}\Bigg{)}^{-1};$
@@
Appendix E How Does FLANDERS Work?
E.1 A Step-by-Step Example
In Figure 2, we depict how FLANDERS computes the anomaly score vector $\bm{s}^{(t)}$ at the generic FL round $t$. Specifically, the server $S$ uses its current MAR($1$) forecasting model $f$ whose best parameters $\hat{\bm{\Omega}}=\{\hat{\bm{A}},\hat{\bm{B}}\}$ are estimated from $l$ previous historical observations of local models received in the previous rounds $t-l,\ldots,t-1$. It applies $f$ to the previously observed matrix $\bm{\bm{\Theta}}_{t-1}$ to get the next predicted matrix of local models $\bm{\hat{\bm{\Theta}}}_{t}$, i.e., $\bm{\hat{\bm{\Theta}}}_{t}=f(\bm{\bm{\Theta}}_{t-1};\hat{\bm{\Omega}})=\hat{\bm{A}}\bm{\bm{\Theta}}_{t-1}\hat{\bm{B}}$. Then, it compares this predicted matrix $\bm{\hat{\bm{\Theta}}}_{t}$ with the actual matrix of local updates received $\bm{\bm{\Theta}}_{t}$. The final anomaly score vector is calculated by measuring the distance $\delta$ between each column of those two matrices, according to Eq. (4) defined in the main paper.
It is worth remarking that, in general, only some clients are selected at every round. In particular, if a client $c_{\text{new}}$ – whether it is honest or malicious – is selected by the server for the first time at round $t$, the MAR forecasting model $f$ will not be able to make any prediction for it due to a cold start problem (i.e., the predicted matrix $\bm{\hat{\bm{\Theta}}}_{t}$ will not contain a column corresponding to $c_{\text{new}}$). In such a case, as already stated in the main submission, we must adopt a fallback strategy. Without any historical information for a client, the most sensible thing to do is to compute the distance $\delta$ between the local model update it sent and the current global model.
Figure 2: Overview of FLANDERS.
[…]
It is worth noting that $\mathcal{C}^{(1)}\cap\mathcal{C}^{(2)}=\{c_{3},c_{9}\}$; the other two clients, $c_{1}$ and $c_{6}$, are considered “cold-start” since they are selected for the first time or, in any case, they do not appear in the historical observations used by MAR for predictions (i.e., in the previous $w=1$ matrices). Thus, we calculate the following anomaly scores, according to Eq. (4) in the main body:
To overcome this problem, FLANDERS replaces the local model marked as suspicious $\bm{\theta}_{3}^{(2)}$ with either one of the previously observed models from the same client that is supposedly legitimate or the current global model. Since, in this example, we assume $c_{3}$ has been malicious from the first round, we use the latter approach. Specifically, we change $\bm{\Theta}_{2}$ with $\bm{\Theta}^{\prime}_{2}=[\bm{\theta}_{1}^{(2)},\underline{\bm{\theta}^{(2)}},\bm{\theta}_{6}^{(2)},\bm{\theta}_{9}^{(2)}]$, where $\bm{\theta}^{(2)}$ substitutes $\bm{\theta}_{3}^{(2)}$. As discussed at the end of Section 5.2 of the main body, this fix allows FLANDERS to work even when a malicious client is picked for two or more rounds consecutively.
Therefore, anomaly scores are updated as follows:
E.2 FLANDERS’ Pseudocode
Algorithm 1 FLANDERS-Server
1:The aggregation function ($\phi$); the number of randomly selected clients at each FL round ($m$); the number of local models to keep as legitimate ($k$); the autoregressive order of MAR ($w$); the number of historical observations used to train MAR ($l$); the number of MAR training iterations ($N$); the number of total FL rounds ($T$).
2:The global model $\bm{\theta}^{(T)}$.
3:procedure FLANDERS-Server($\phi$, $m$, $k$, $w$, $l$, $N$, $T$)
4:     $\bm{\theta}^{(1)}\leftarrow$ A randomly initialized model
5:     for each $t\in\{1,2,\ldots T\}$ do
6:         $\mathcal{C}^{(t)}\leftarrow$ sample a subset of $m$ clients from $\mathcal{C}$
7:         Send the global model $\bm{\theta}^{(t)}$ to every $c\in\mathcal{C}^{(t)}$
8:         $\bm{\Theta}_{t}\leftarrow[\bm{\theta}^{(t)}_{1},\ldots,\bm{\theta}^{(t)}_{m}]$triangleright Receive the $m$ local models trained by each $c\in\mathcal{C}^{(t)}$
9:         if $t==1$ then
10:              $\mathcal{C}^{(t)}_{*}\leftarrow\texttt{{Fallback}}(\bm{\Theta}_{t})$triangleright At the very first round, use the designated fallback aggregation strategy (e.g., FedAvg)
11:         else
12:              $\mathcal{C}^{(t)}_{*}\leftarrow\texttt{{FLANDERS-Filter}}(\bm{\Theta}_{t-l:t-1},\mathcal{C}^{(t)},k,w,N)$triangleright Returns the set of clients classified as legitimate
13:         end if
14:         $\bm{\theta}^{(t+1)}\leftarrow\phi(\{\bm{\theta}_{c}^{(t)}~{}|~{}c\in\mathcal{C}_{*}^{(t)}\})$
15:     end for
16:end procedure
Algorithm 2 FLANDERS-Filter
1:The tensor containing the local models of selected clients observed from $l$ past FL rounds ($\bm{\Theta}_{t-l:t-1}$); the clients selected at round $t$ ($\mathcal{C}^{(t)}$); the autoregressive order of MAR ($w$); the number of local models to keep as legitimate ($k$); the number of MAR training iterations ($N$).
2:The set of $k$ clients classified as legitimate in round $t$ $\mathcal{C}^{(t)}_{*}$.
3:procedure FLANDERS-Filter($\bm{\Theta}_{t-l:t-1},\mathcal{C}^{(t)},k,w,N$)
4:     $\hat{\bm{\Theta}}_{t}\leftarrow\texttt{MAR}(\bm{\Theta}_{t-l:t-1},w,N)$triangleright MAR($w$) estimation via ALS training ($N$ iterations)
5:     $\bm{s}^{(t)}\leftarrow\delta(\bm{\Theta}_{t},\hat{\bm{\Theta}}_{t})$triangleright Anomaly score vector
6:     $\mathcal{C}^{(t)}_{*}\leftarrow\{c\in\mathcal{C}^{(t)}~{}|~{}s_{c}^{(t)}\leq s_{k}^{(t)}\}$triangleright $s^{(t)}_{k}$ is the $k$-th smallest anomaly score
7:     return $\mathcal{C}^{(t)}_{*}$
8:end procedure
Appendix F Datasets, Tasks, and FL Models

### v3

@@ VIII Conclusion
In the future, we will address the primary limitations of this work, as discussed in Section VII.
@@
The (anonymous) GitHub repository with the code and the data to replicate the results discussed in this work is accessible at the following link:
https://anonymous.4open.science/r/flanders_exp-7EEB
IX A Note on the Terminology Used
@@
X The Impact of Malicious Clients at each FL Round
@@
XI Matrix Autoregressive Model (MAR)
@@
$\begin{split}\ell(\bm{\widetilde{\Omega}};{\bm{\bm{\Theta}}}_{t})&=||{\bm{\bm{\Theta}}}_{t}-\widetilde{{\bm{\bm{\Theta}}}}_{t}||^{2}_{\text{F}}\\ &=||{\bm{\bm{\Theta}}}_{t}-f(\bm{\Theta}_{t-1};\bm{\widetilde{\Omega}})||^{2}_{\text{F}}\\ &=||{\bm{\bm{\Theta}}}_{t}-\widetilde{{\bm{A}}}{\bm{\bm{\Theta}}}_{t-1}\widetilde{{\bm{B}}}||^{2}_{\text{F}}\\ &=\ell(\widetilde{{\bm{A}}},\widetilde{{\bm{B}}};{\bm{\bm{\Theta}}}_{t}),\end{split}$ (7)
@@
$\begin{split}\hat{{\bm{\Omega}}}=\hat{\bm{A}},\hat{\bm{B}}&=\text{arg\,min}_{\widetilde{{\bm{A}}},\widetilde{{\bm{B}}}}\Big{\{}\mathcal{L}(\widetilde{{\bm{A}}},\widetilde{{\bm{B}}};{\bm{\bm{\Theta}}}_{t},l)\Big{\}}\\ &=\text{arg\,min}_{\widetilde{{\bm{A}}},\widetilde{{\bm{B}}}}\Big{\{}\sum_{j=0}^{l-1}||{\bm{\bm{\Theta}}}_{t-j}-\widetilde{{\bm{A}}}{\bm{\bm{\Theta}}}_{t-j-1}\widetilde{{\bm{B}}}||^{2}_{\text{F}}\Big{\}}.\end{split}$ (9)
@@
$\small\begin{split}\frac{\partial{\mathcal{L}({\bm{A}},{\bm{B}};{\bm{\bm{\Theta}}}_{t},l)}}{\partial{{\bm{A}}}}&=\frac{\partial{\sum_{j=0}^{l-1}||{\bm{\bm{\Theta}}}_{t-j}-{\bm{A}}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}||^{2}_{\text{F}}}}{\partial{{\bm{A}}}}=\\ &=-2\sum_{j=0}^{l-1}({\bm{\bm{\Theta}}}_{t-j}-{\bm{A}}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}){\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-1}=\\ &=-2\Bigg{[}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\\ &-{\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\Bigg{]}.\end{split}$ (11)
@@
$\footnotesize\begin{split}-2\Bigg{[}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}-{\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\Bigg{]}=0\end{split}$
@@
$\footnotesize\begin{split}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}-{\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}=0.\end{split}$
@@
$\footnotesize\begin{split}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}-{\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}=0\end{split}$ (12)
$\footnotesize\begin{split}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}={\bm{A}}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\end{split}$ (13)
$\footnotesize\begin{split}{\bm{A}}=\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}{\bm{B}}^{T}{\bm{B}}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}^{-1}.\end{split}$ (14)
@@
$\small\begin{split}{\bm{B}}=\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}^{T}_{t-j}{\bm{A}}{\bm{\bm{\Theta}}}_{t-j-1}\Bigg{)}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}^{T}_{t-j-1}{\bm{A}}^{T}{\bm{A}}{\bm{\bm{\Theta}}}_{t-j-1}\Bigg{)}^{-1}.\end{split}$ (16)
@@
$\scriptsize\begin{split}{\bm{A}}^{(i+1)}=\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j}{\bm{B}}^{(i)}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}_{t-j-1}({\bm{B}}^{(i)})^{T}{\bm{B}}^{(i)}{\bm{\bm{\Theta}}}^{T}_{t-j-1}\Bigg{)}^{-1};\end{split}$
$\scriptsize\begin{split}{\bm{B}}^{(i+1)}=\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}^{T}_{t-j}{\bm{A}}^{(i+1)}{\bm{\bm{\Theta}}}_{t-j-1}\Bigg{)}\Bigg{(}\sum_{j=0}^{l-1}{\bm{\bm{\Theta}}}^{T}_{t-j-1}({\bm{A}}^{(i+1)})^{T}{\bm{A}}^{(i+1)}{\bm{\bm{\Theta}}}_{t-j-1}\Bigg{)}^{-1};\end{split}$
@@
XII Datasets, Tasks, and FL Models